Consider the function φ (x,y)-ex + e. Consider the point (x,y)= (1,1). Dyalculate the exact values ofipor and 0φ。y at this point b) Ue first-order forward differences, with Ax Ay 0., to calculate approximate valueoföx and öpy at point (1, 1). Calculate the percentage difference when compared with the exact values from part (a). (c) Use first-order rearward differences, with Ar = 0.1, to calculate approximate values of 6fax and öply at point (1, Ealculate the percentage difference when mpared with the exact values fpem part (a) -((d) Uje second-order central differences, with Ar-Av-0.1, to calculate approx- ー rifate values for ablar and ahoy at point (1, 1). Calculate the percentage difference when compared with the exact value from part (a).
Answer #1

오을부(1,1,-lenge)-(e+e)- =2-8S88 o 밖,, ngtete): - (ete 8588 ce- d8588-978f83 -718 3 O Cee 25869 1C,I) ete こ3-586 8 2-718 3-2-ах. 山 lenge)-( ·リー or 흠 3x0.1 gnance = ce2:7228-2-7133to 166 y. 9. 2.7183